Show that the expression for the electric field of a charged disk showing in the equation below reduces to that of a point for Z>>R, where all symbols have their usual meanings.



σ/2εo[ 1= z/√z^2+R^2]

1
Expert's answer
2021-11-25T11:51:32-0500

The electric field of a charged disk


"E=\frac{\sigma}{2\epsilon_0}(1-\frac{z}{\sqrt{z^2+R^2}})=\frac{\sigma}{2\epsilon_0}(1-\frac{1}{\sqrt{1+\frac{R^2}{z^2}}})"


For "z\gg R"


"\frac{1}{\sqrt{1+\frac{R^2}{z^2}}}\approx1-\frac{R^2}{2z^2}" . So, we have


"E=\frac{\sigma}{2\epsilon_0}(1-\frac{1}{\sqrt{1+\frac{R^2}{z^2}}})\approx\frac{\sigma R^2}{4\epsilon_0h^2}=\frac{q}{4\pi\epsilon_0z^2}" as for the point charge field.